OpenStudy (anonymous):

Solve 11.3^2x – 1 = 15.7

OpenStudy (anonymous):

11.3^2x=16.7 take log on both sides and solve

OpenStudy (anonymous):

you forgot the - 1..

OpenStudy (anonymous):

no added 11.3^2x=15.7+1 11.3^2x=16.7

OpenStudy (anonymous):

x ≈ 1.12 x ≈ 0.57 x ≈ 0.94 x ≈ 1.07 these are my options..

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

x approximately equal to 0.57

OpenStudy (anonymous):

right?